Shift report 20011016 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Tue Oct 16 2001 - 00:05:38 EDT

  • Next message: Dieter Rohrich: "Re: BRAHMS NIM Paper "final" revision"

    ------------------------------------------------------------
    Date: Tue Oct 16 00:05:37 2001, Shift: 16:00-24:00
    Supervisor: JH Lee, with: 
    Log book#10 pages 13 - 15
    
    Shift Summary:
    ------------------------------------------------------------
    Beam:  
     ~5 Hours Store(starting ZDC rates at 400/sec)
    
    Data: 
     Physics Runs: #5161,#5162 (0.64M Trigger1)
     at MRS 45deg 4B and FS 4/4deg 1/5B
    
    Changes:
     - New HV settings for C1 
       "now all PMT show 1 p.e. 40 or more
        channels above pedestals" (RD)
     - T5 Threshold Voltage set to 3.2 -> 3.0 by 
     a request from Zbigniew
    
    Problems:
     - BBL#37 HV found to be off->Reset: 
       This has happened before. Might need to
       be checked.
     - TOFW Top92 as reported by previous 
       shifters. 
      
     
           
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Tue Oct 16 2001 - 00:06:03 EDT